Please help me solve this word problem:
A drug store parking lot has space for 1000 cars. 2/5 of the spaces are for compact cars. On Tuesday, there were 200 compact cars and some standard size cars in the parking lot. The parking lot was 3/4 full. How many standard size cars were in the parking lot?

----Thank you =)

Answer by ankor@dixie-net.com(6693)

A drug store parking lot has space for 1000 cars. 2/5 of the spaces are for compact cars.
* 1000 = 400 spaces available for compact cars
then
1000-400 = 600 spaces available for standard cars
:
On Tuesday, there were 200 compact cars and some standard size cars in the parking lot.
Let x = no. of standard cars
:
The parking lot was 3/4 full. How many standard size cars were in the parking lot?
:
x + 200 = *1000
x + 200 = 750
x = 750 - 200
x = 550 standard cars (600 spaces available)
;
:
We can check solution this way
200 + 550 = 750 which 3/4 of 1000